Please note: You are viewing the unstyled version of this website. Either your browser does not support CSS (cascading style sheets) or it has been disabled. Skip navigation.

May 29, 2014 - ESP Update

   Print Change text to small (default) Change text to medium Change text to large

Here's the latest ESP results as of yesterday (May 28, 2014).

-->ESP-1 (don) deployed near NERACOOS B detected a return of the A. fundyense population in the 300 cells/L range after the spike in the population last week.

-->ESP-2 (jake) in Casco Bay also showed a slight uptick to ~200 cells/Liter after the spike last week.

-->ESP-3 (dennis) deployed offshore of Pemaquid point, had some technical difficulties yesterday. Thanks to the engineers at MBARI, ESPdennis' run was recovered. No positive A. fundyense signals were observed on the compromised "hab" array (not plotted). However, the MUD2 probe (1 of 4 Pseudo-nitzschia probes on the array) did show positive hits.

Bruce Keafer

Last updated: July 1, 2019

whoi logo

Copyright ©2007 Woods Hole Oceanographic Institution, All Rights Reserved, Privacy Policy.
Problems or questions about the site, please contact